### v1

# Universal low-depth two-unitary design of programmable photonic circuits

Abstract
The development of large-scale, programmable photonic circuits capable of performing generic matrix-vector multiplication is essential for both classical and quantum information processing. However, this goal is hindered by high losses, hardware errors, and difficulties in programmability. We propose an enhanced architecture for programmable photonic circuits that minimizes circuit depth and offers analytical programmability, properties that have not been simultaneously achieved in previous circuit designs. Our proposal exploits a previously overlooked representation of general nonunitary matrices as sums of two unitaries. Furthermore, similar to the traditional SVD-based circuits, the circuits in our unitary-sum-based architecture inherit the advantages of the constituent unitary circuits. Overall, our proposal provides a significantly improved solution for matrix-vector multiplication compared to the established approaches.
Introduction
Programmable photonic circuits are increasingly used as an energy-efficient solution for classical information processing [1, 2, 3] and play a crucial role in optical quantum computing [4, 5, 6, 7]. However, as analog devices, photonic circuits face challenges such as errors, losses, and limited programmability that hinder their scalability and prevent them from competing with traditional digital electronic methods.
The core mathematical operation that underpins the computational power of programmable multimode photonics is matrix-vector multiplication (MVM). This operation is vital for data transformation [8, 9, 10], weight adjustment, and learning processes in photonic neural networks [1]. Recent advancements in photonic architectures have largely focused on programmable interferometers that perform multiplications with unitary transfer matrices. Several unitary designs for programmable photonic circuits have been proposed and widely adopted by researchers. Notably, the designs by Reck et al.[11] and Clements et al.[12] stand out, as they benefit from an analytical procedure that calculates the phase-shift values needed for a given target matrix, assuming the circuits are error-free. Additionally, methods for error correction in unitary programmable circuits have been suggested [13], along with novel circuit designs that maintain universality even in the presence of high error rates [14, 15, 16].
However, many information processing tasks require MVM involving a broader range of matrices beyond the unitary group. This is common in classical photonic neural networks [3], iterative solvers [2], and quantum graph problem solvers [7]. Two main approaches have been proposed to implement photonic multiplication by nonunitary matrices. The first is the well-established singular value decomposition (SVD) method, which uses a sequence of a programmable unitary multimode circuit, mode-wise amplitude modulation, and another unitary circuit [17]. The advantage of this approach is its straightforward programmability, inherited from the analytically programmable unitary circuits.
The second approach involves embedding the target nonunitary matrices into larger unitary matrices [18, 19, 16]. These methods typically require shorter circuit depths compared to SVD-based circuits. For example, Tang et al. [18] proposed embedding target nonunitary matrices into unitary circuits of half the depth of the corresponding SVD-based circuits. More recently, we proposed an even more compact architecture that is compatible with planar integrated photonics [19]. However, both approaches sacrifice the ease of analytical programmability, making their practical use challenging due to the lack of effective programming algorithms and in situ training methods. This, in turn, can spoil all the computational advantage of photonic MVM.
In this work, we propose a new programmable photonic circuit architecture that significantly improves both the SVD-based circuits [17] and our recent low-depth architecture [19]. The proposed architecture is based on representing the target nonunitary matrices as the sum of two unitary matrices. Our approach combines the advantages of both the SVD-based and embedding-into-unitary methods. Specifically, the programmable circuits derived from our architecture are both low-depth and analytically programmable.
I Low-depth circuit
We are interested in generic MVM operation, in which the input vector of field amplitudes $\boldsymbol{a}^{(in)}=(a^{(in)}_{1},\ldots,a^{(in)}_{N})^{\text{T}}$ is multiplied by a target $M\times{}N$ complex-valued matrix $W$:
$\boldsymbol{a}^{(out)}=W\boldsymbol{a}^{(in)},$ (1)
in the course of their propagation through a programmed multiport linear-optical circuit to obtain the result $\boldsymbol{a}^{(out)}=(a^{(out)}_{1},\ldots,a^{(out)}_{M})^{\text{T}}$ at its output. In the following, we assume square matrices $W$ with $N=M$, which will not undermine the generality of our results. The general task of performing MVM by optics boils down to constructing a programmable multiport circuit that can implement a class of transfer matrices of interest $W$.
In this work, we are interested in the universal MVM — the most general form of MVM, in which the circuit is required to implement arbitrary complex-valued matrices $W$, which can be parametrized by $2N^{2}$ real parameters. To this date, the most often exploited approach to accomplish this goal has been using the SVD-based circuits, depicted in Fig. 1. The SVD represents any non-unitary complex-valued matrix $W$ in the form of the product of three matrices:
$W=U\Sigma{}V^{\dagger},$ (2)
each of which is implemented by a distinct photonic unit, as shown in Fig. 1. In (2), $U$ and $V$ are unitary matrices and $\Sigma=\text{diag}(\sigma_{1},\ldots,\sigma_{N})$ is the diagonal matrix consisting of the real singular values $\sigma_{j}$ of the matrix $W$. Strictly speaking, for arbitrary matrices $W$ the singular values $\sigma_{j}$ are unbounded, however, the linear optical implementation of the SVD imposes the constraint $|\sigma_{j}|\leq{}1$. Fortunately, this does not diminish the applicability of linear optical implementations, since the target matrices $W$ can be rescaled to meet the constraint.
Figure 1: The SVD-based programmable circuit for the universal MVM operation.
I.0.1 Proposal
The circuit we propose in this work is based on the simple observation that the product (2) can be rewritten as a sum of two unitary matrices. To show this, we rewrite the diagonal matrix $\Sigma$ as follows:
$\Sigma=\frac{1}{2}(D+D^{\ast}),$ (3)
where $D=\text{diag}(e^{i\psi_{1}},\ldots,e^{i\psi_{N}})$ is the diagonal unitary matrix with $e^{i\psi_{j}}=\sigma_{j}+i\sqrt{1-\sigma_{j}^{2}}$ ($j=\overline{1,N}$). Then, substituting (3) into (2), we come to the expression:
$W=\frac{1}{2}(U{}D{}V^{\dagger}+U{}D^{\ast}V^{\dagger}).$ (4)
Figure 2: The programmable circuit implementing the sum of two unitaries $U^{(1)}$ and $U^{(2)}$ proposed to perform multiplication by $N\times{}N$ nonunitary matrices. Here, the BSs before and after unitary circuits are balanced.
The optical scheme performing MVM according to (4) is shown in Fig. 2. The main part of the scheme is formed by two $N$-mode unitary circuits, $U^{(1)}=U{}D{}V^{\dagger}$ and $U^{(2)}=U{}D^{\ast}V^{\dagger}$, acting in parallel rather than in sequence, as in the SVD-based circuits shown in Fig. 1. The MVM multiplication proceeds by first splitting the field vector to be multiplied $\boldsymbol{a}^{(in)}$ on a set of $N$ balanced beam-splitters BS, so that each its component is equally distributed among the unitary circuits. Without loss of generality, we take the beam-splitter transfer matrix as
$U_{\text{BS}}=\left(\begin{array}[]{cc}\sqrt{R}&i\sqrt{1-R}\\ i\sqrt{1-R}&\sqrt{R}\end{array}\right),$ (5)
where $R$ is the power reflectivity of the BS that should ideally be balanced (i.e. $R=R_{0}=1/2$). In the following, we write $R=\cos^{2}(\pi/4+\alpha)$, where $\alpha$ quantifies the deviation of the BS from the balanced operation. After multiplication by the unitary circuits the amplitudes are combined on yet another set of balanced beam-splitter BS to produce $2N$ amplitudes $a^{(out)}_{1},\ldots,a^{(out)}_{2N}$. Then, the MVM result (1) is carried by the $N$ modes $a^{(out)}_{1},\ldots,a^{(out)}_{N}$.
The circuit exploiting two parallel programmable unitaries has the obvious advantage of having two times lower depth than the SVD-based counterpart, which results in proportionally lower losses. At the same time, our circuit architecture retains the advantage of being easily reprogrammed according to a given target matrix without resorting to the computationally-intensive numerical optimization inherent to other low-depth circuits [19, 18].
I.0.2 Tolerance to hardware errors
Taking into account various error-tolerant designs for universal unitaries, such as circuits [14, 15, 16], the non-unitary circuit under investigation can also be made error-tolerant, provided that the errors in the BSs connecting two unitaries are corrected. This can be achieved by replacing each of the $2N$ BSs with a programmable Mach-Zehnder interferometer, introducing a linear overhead with respect to $N$, which is relatively small compared to the quadratic overhead from the programmable elements in the unitaries. However, in our study, we focus on the impact of hardware errors while assuming that these BSs remain static, but the transfer matrices $U^{(1)}$ and $U^{(2)}$ are unaffected by errors. We do not assume a specific error-tolerant photonic architecture for the unitary circuits; instead, we treat them as parameterized matrices. Specifically, we parametrize the unitaries $U^{(1)}$ and $U^{(2)}$ by $2N^{2}$ parameters of the non-unitary matrix $K=\frac{U^{(1)}+U^{(2)}}{2}$ with singular values not exceeding $1$. It should be noted that if no hardware errors are present, $K=W$, however, in the general erroneous case $K\neq W$.
Thus, the only faulty elements are the BSs (5), with the values of $\alpha$’ representing the corresponding errors. Two limiting cases of error models were considered: i) correlated errors, where all BSs have the same error $\alpha$, and ii) independent random errors, where the error $\alpha_{j}$ for each BS follows a normal distribution $\sim\exp(-\alpha_{j}^{2}/2\sigma^{2})$, with $\sigma$ representing the degree of randomness in the errors.
Figure 3: Effect of beam-splitter errors on the quality of photonic MVM operations. The distributions of $\mathrm{RMSE}$ and $\mathrm{RMSE_{s}}$ are shown for i) coherent errors as a function of the BS power reflectivity $R$ at $N=10$ (a) and $N=20$ (b), and ii) independently sampled random errors as a function of the distribution width $\sigma$ at $N=10$ (c) and $N=20$ (d). Each $\mathrm{RMSE}$/$\mathrm{RMSE_{s}}$ distribution is based on $100$ randomly sampled target matrices. The distributions of $|s|$ corresponding to $\mathrm{RMSE_{s}}$ are also plotted in the figures. For the distributions corresponding to independently sampled random errors the minium, maximum and median bars are plotted.
To assess how accurately the actual circuit implements the MVM under BS errors, we compare the target matrices $W^{(0)}$ with the corresponding transfer matrices $W$ realized by the circuit. For this, we use the root mean square error (RMSE):
$\text{RMSE}(W^{(0)},W)=\frac{1}{N}\sqrt{\sum_{i,j=1}^{N}|W^{(0)}_{ij}-W_{ij}|^{2}},$ (6)
where $W_{ij}$ and $W_{ij}^{(0)}$ are the elements of the matrices. Additionally, we use a more relaxed measure of matrix closeness, allowing $W^{(0)}$ and $W$ to coincide up to a global scaling factor $s$. The corresponding measure of closeness is $\mathrm{RMSE_{s}}(W^{(0)},W)=\mathrm{RMSE}(sW^{(0)},W)/|s|$ [19]. Smaller values of $|s|$ are associated with larger losses, so they are penalized by $|s|$ in the denominator in $\mathrm{RMSE_{s}}$. To determine the parameters of the unitary circuits that implement a target non-unitary matrix $W^{(0)}$ with the lowest possible $\mathrm{RMSE}$/$\mathrm{RMSE_{s}}$, we employed the L-BFGS optimization algorithm [20]. In optimizing $\mathrm{RMSE_{s}}$, we used the method from our previous work [19] to find the scaling factor $s$ that minimizes MVM error while ensuring that its absolute value is maximized.
We used SVD (2) to generate a target matrix $W^{(0)}$, wherein the unitary complex-valued matrices $U$ and $V$ were drawn from the Haar random distribution using the method based on the QR-decomposition of random matrices from the Ginibre ensemble [21]. The diagonal matrix $\Sigma$ was filled with independently generated values from a uniform distribution in the $[0,1]$ range, which are then rescaled in such a way that the maximum value of $\sigma_{j}$ is equal to $1$.
Fig. 3 shows the results of the simulation. In Figs. 3a and 3b the $\mathrm{RMSE}$ and $\mathrm{RMSE_{s}}$ distributions are plotted as a function of the power reflectivity for the circuit sized of $N=10$ and $N=20$, respectively. As seen in the figures, even a small deviation of $R$ from its balanced value of $1/2$ degrades the quality of the matrix operation. However, when the transfer matrix is allowed to scale by the global factor $s$, the coherent errors can be perfectly corrected by the unitary circuits and measure the quality of the MVM by $\mathrm{RMSE_{s}}$. Notice that the absolute value of the scaling factor required to correct coherent errors monothonically decreases with increasing value of error.
Fig. 3c and Fig. 3d suggest different behavior of the MVM quality, as quantified by both unscaled and scaled measures. However, the implementation of scaled matrices, quantified by $\mathrm{RMSE_{s}}$, is much better than that of unscaled matrices, quantified by $\mathrm{RMSE}$. It should be noted that this correction is attained by scaling factors that differ from $1$ by small values.
In summary, we propose an architecture for programmable photonic circuits capable of performing MVM operations, inspired by the decomposition of nonunitary matrices into a sum of two unitaries – an approach previously used in quantum computing theory [22] and, to our knowledge, never applied in the context of photonic MVMs. This architecture offers the advantage of having half the circuit depth compared to SVD-based interferometer circuits, which are the standard approach to performing MVM with programmable photonics today. Moreover, using the proposed programmable circuits is a more efficient choice than the recently investigated low-depth circuits [19, 18], as our circuits can be programmed analytically while maintaining the same circuit depth. We believe that our results will be valuable for various applications in photonic information processing.
II Acknowledgments
S.A.F. is grateful to the Russian Foundation for the Advancement of Theoretical Physics and Mathematics (BASIS) (Project №23-2-10-15-1).